OpenStudy (anonymous):

I am confuse in composition of function lemme show you a example tell me is it a composition of function f(x)=x2+1 g=x,h=x2+1 g(h(x))=x2+1 what is this is this a composition of fuction? or something else

OpenStudy (anonymous):

Yes,this is a composition of a function!! Your question says that g(x)=x and h(x)= x^2 + 1.Therefore, g(h(x)) = g(x^2 + 1) [as h(x)=x^2 + 1] = x^2 + 1 [as g(x)=x]

OpenStudy (anonymous):

ok thnk you

OpenStudy (anonymous):

My pleasure!!

OpenStudy (anonymous):

can you tell me about reflection of function